OpenStudy (tjhub00):

What is the value of x in the equation −x = 3 − 4x + 6? −9 −3 3 9

OpenStudy (anonymous):

Let's solve your equation step-by-step. −x=3−4x+6 Step 1: Simplify both sides of the equation. −x=3−4x+6 Simplify: (Show steps) −x=−4x+9 Step 2: Add 4x to both sides. −x+4x=−4x+9+4x 3x=9 Step 3: Divide both sides by 3. \[\frac{ 3x }{ 3 }\] \[\frac{ 9 }{ 3 }\] x = ? What is 9 divided by 3 @tjhub00?

OpenStudy (tjhub00):

thank you! :)